# Data-Driven Linearization of Dynamical Systems  [^1] To appear in Nonlinear Dynamics

### v1

@@ Appendix B Proof of Theorem 1 and technical remarks
In most applications the dynamical system will have higher degree of smoothness, i.e., we will have $\mathbf{f}\in C^{r}$ for some $r\in\mathbb{N}^{+}\cup\left\{\infty,a\right\}$, with $C^{a}$ referring to the space of analytic functions. Then, under further nonresonance conditions of the spectrum of $\mathbf{A}$, the spectral submanifold $\mathcal{W}(E)$ will also be of class $C^{r}$, as we discuss in Section 4. In those cases, for the approximate topological equivalence holds in the statement of Theorem 1 hold on the full domain of attraction of the $\mathbf{\mathbf{x}}=\mathbf{0}$ fixed point within $\mathcal{W}(E)$, as one deduced from the linearization-in-the-large results of Lan and Mezic [35], Eldering et al. [17]. [^6] Note that these results require at least $C^{2}$ differentiability for the dynamical system. Consequently, one cannot use them within $\mathcal{W}(E)$ under the general assumptions of Theorem 1, which only guarantee $C^{1}$ differentiability for the reduced flow in $\mathcal{W}(E)$. In Theorem 1, our objective was to give a minimal set of conditions under which DMD can be justified as an approximate, leading-order, $d$-dimensional reduced model for the original nonlinear system. These minimal conditions only require hyperbolicity, i.e., robustness of the spectrum under perturbations, without insisting on the lack of resonances in the spectrum. This is advantageous in data-driven applications in which details of the spectrum are generally not known.
@@ Appendix D Proof of Theorem 4
with $\mathbf{h}\in C^{r}$. Additionally, a strengthening of Sternberg’s linearization theorem by Eldering et al. [17] [^7] This is based on the approach of Lan and Mezic [35] in the setting of the $C^{1}$ linearization theorem by Hartman [23]. implies that the transformation in eq. (106) is unique and globally defined on the full domain of attraction of the fixed point at the origin. As pointed out by Haller et al. [22], this implies the existence of a unique $d$-dimensional, normally attracting spectral submanifold $\mathcal{W}(E)\in C^{\infty}$ for system (11) whose internal dynamics govern the longer-term behavior of all trajectories near the origin. [^8] Cabré et al. [10] guarantees the existence and uniqueness of $\mathcal{W}(E)\in C^{\infty}$ under weaker nonresonance conditions that allow resonances inside $E$ but no resonances of the form (105) with $\lambda_{k}\in\mathrm{Spect}\,\left(A|_{F}\right)$ and $\lambda_{j}\in\mathrm{Spect}\,\left(\mathbf{A}|_{E}\right).$ We do not rely on these results here, as the next step in our construction is local linearization within $\mathcal{W}(E)$, which cannot be carried out in the presence of resonances inside $E$.
@@
$Q=\mathrm{Int}\left[\frac{\max_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}{\min_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}\right]+1,$
@@
We have seen that $D_{\boldsymbol{\xi}}\boldsymbol{\varphi}(\mathbf{0})$ is nonsingular and hence the linear part of the nonlinear ODE in (112) has the same eigenvalues as $\boldsymbol{\Lambda}_{E}$ does. $\left(\mathbf{T}^{-1}\mathbf{A}\mathbf{T}\right)|_{E}$. Consequently, the origin of system (113) is a hyperbolic fixed point. If the nonresonance conditions (36) hold for the full spectrum of $\mathbf{A}$, then they also hold for any subset of this spectrum, i.e., for the eigenvalues of $\boldsymbol{\Lambda}_{E}$ as well. Consequently, from the linearization theorems of Sternberg [52] and Eldering et al. [17], we can also deduce a similar near-identity, $C^{r}$ linearizing diffeomorphism

### v2

@@ Appendix B Proof of Theorem 1 and technical remarks
In most applications the dynamical system will have higher degree of smoothness, i.e., we will have $\mathbf{f}\in C^{r}$ for some $r\in\mathbb{N}^{+}\cup\left\{\infty,a\right\}$, with $C^{a}$ referring to the space of analytic functions. Then, under further nonresonance conditions of the spectrum of $\mathbf{A}$, the spectral submanifold $\mathcal{W}(E)$ will also be of class $C^{r}$, as we discuss in Section 4. In those cases, for the approximate topological equivalence holds in the statement of Theorem 1 hold on the full domain of attraction of the $\mathbf{\mathbf{x}}=\mathbf{0}$ fixed point within $\mathcal{W}(E)$, as one deduces from the linearization-in-the-large results of Lan and Mezic [35] and Kvalheim and Revzen [34]. [^5] Note that these global linearization results rely on the existence of a $C^{1}$ local linearization, which generally only exists for class $C^{2}$ dynamical systems. Consequently, one cannot use the available global linearization results within $\mathcal{W}(E)$ under the general assumptions of Theorem 1, which only guarantee $C^{1}$ differentiability for the reduced flow in $\mathcal{W}(E)$. In Theorem 1, our objective was to give a minimal set of conditions under which DMD can be justified as an approximate, leading-order, $d$-dimensional reduced model for the original nonlinear system. These minimal conditions only require hyperbolicity, i.e., robustness of the spectrum under perturbations, without insisting on the lack of resonances in the spectrum. This is advantageous in data-driven applications in which details of the spectrum are generally not known.
@@ Appendix D Proof of Theorem 4
with $\mathbf{h}\in C^{r}$. Additionally, a strengthening of Sternberg’s linearization theorem by Kvalheim and Revzen [34] [^6] This generalizes the results of Lan and Mezic [35] in the setting of the $C^{1}$ linearization theorem by Hartman [23]. implies that the transformation in eq. (106) is unique and globally defined on the full domain of attraction of the fixed point at the origin. As pointed out by Haller et al. [22], this implies the existence of a unique $d$-dimensional, normally attracting spectral submanifold $\mathcal{W}(E)\in C^{\infty}$ for system (11) whose internal dynamics govern the longer-term behavior of all trajectories near the origin. [^7] Cabré et al. [10] guarantees the existence and uniqueness of $\mathcal{W}(E)\in C^{\infty}$ under weaker nonresonance conditions that allow resonances inside $E$ but no resonances of the form (105) with $\lambda_{k}\in\mathrm{Spect}\,\left(A|_{F}\right)$ and $\lambda_{j}\in\mathrm{Spect}\,\left(\mathbf{A}|_{E}\right).$ We do not rely on these results here, as the next step in our construction is local linearization within $\mathcal{W}(E)$, which cannot be carried out in the presence of resonances inside $E$.
@@
$Q=\left\lfloor\frac{\max_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}{\min_{i}\left|\mathrm{Re}\,\lambda_{i}\right|}\right\rfloor+1,$
@@
We have seen that $D_{\boldsymbol{\xi}}\boldsymbol{\varphi}(\mathbf{0})$ is nonsingular and hence the linear part of the nonlinear ODE in (112) has the same eigenvalues as $\boldsymbol{\Lambda}_{E}$ does. $\left(\mathbf{T}^{-1}\mathbf{A}\mathbf{T}\right)|_{E}$. Consequently, the origin of system (113) is a hyperbolic fixed point. If the nonresonance conditions (36) hold for the full spectrum of $\mathbf{A}$, then they also hold for any subset of this spectrum, i.e., for the eigenvalues of $\boldsymbol{\Lambda}_{E}$ as well. Consequently, from the linearization theorems of Sternberg [52] and Kvalheim and Revzen [34], we can also deduce a similar near-identity, $C^{r}$ linearizing diffeomorphism